OpenStudy (jarp0120):

IF AN EDGE OF A CUBE IS INCREASED BY 33%, BY HOW MUCH IS THE TOTAL SURFACE AREA INCREASED? PLEASE THANK YOU :)

OpenStudy (anonymous):

\(\sf Surface\ area =6(edge)^2=6a^2\) edge increase by 33% \(\sf Surface\ area' =6(1.33a)^2= 6(1.76)a^2\) \(\sf Surface\ area' - Surface\ area = 6(1.76)a^2-6a^2=4.56%\) not sure though

OpenStudy (anonymous):

*oops my bad! last line should be: \(\sf Surface\ area' - Surface\ area = 6(1.76)a^2-6a^2=1.76-1=76 %\)

OpenStudy (anonymous):

76%

OpenStudy (jarp0120):

thank you

OpenStudy (anonymous):

not 100% sure if i'm right though

OpenStudy (jarp0120):

ok thank you again:)

OpenStudy (anonymous):

np :)